OpenStudy (anonymous):

I'm really confused on how to help my daughter with this question what three factors were used to make this Game Product board 4, 6,9, 14,? and 49

OpenStudy (anonymous):

you can divide all of these numbers using only 2, 3 and 7

OpenStudy (anonymous):

4 = 2*2 6 = 2*3 9 = 3*3 14 = 2*7 49 = 7*7

OpenStudy (anonymous):

But how do I fiqure out the ?

OpenStudy (anonymous):

the only one missing is 21, 3*7

OpenStudy (anonymous):

Thanks.......I have another one it has four factors 9,15,18, ?, 21, ?, 30,35,?,36,42,49